Lean Accounting

Com-Tel Inc. manufactures and assembles two models of smartphones-the Tiger Model and the Lion Model. The process consists of a lean cell for each product. The data that follow concern only the Lion Model lean cell.

For the year, Com-Tel Inc. budgeted these costs for the Lion Model production cell:

Conversion Cost Categories Budget
Labor $63,400
Supplies 24,000
Utilities 8,600
   Total $96,000

Com-Tel plans 2,000 hours of production for the Lion Model cell for the year. The materials cost is $41 per unit. Each assembly requires 24 minutes of cell assembly time. There was no May 1 inventory for either Raw and In Process Inventory or Finished Goods Inventory.

The following summary events took place in the Lion Model cell during May:

  1. Electronic parts were purchased to produce 6,800 Lion Model assemblies in May.
  2. Conversion costs were applied for 6,450 units of production in May.
  3. 6,320 units were completed and transferred to finished goods in May.
  4. 6,130 units were shipped to customers at a price of $192 per unit.

If required, round your answers to the nearest cent.

Required:

1. Determine the budgeted cell conversion cost per hour.
$ per hour

2. Determine the budgeted cell conversion cost per unit.
$ per unit

3. Journalize the summary transactions (a) through (d).

a.
b.
c.
d. Sale
d. Cost

4. Determine the ending balance in Raw and In Process Inventory and Finished Goods Inventory.

Raw and In Process Inventory $
Finished Goods Inventory $

5. Lean accounting is different from traditional accounting because it is more   and uses   control. As a result, the number of transactions are  . In many lean operations, purchased materials are charged to a  . Direct labor is frequently  . Often, nonfinancial performance measures, such as  , are used to monitor performance.
Answer #1

Answer:

1)Conversion cost per hour = $96,000 /2000 = $48 per hour
2)budgeted cell conversion cost per unit. = Manufacturing Time × Cell Conversion Cost Rate
= 24/60×$48 = $19.20 per unit
3)
Sr. No Accounts Titles Debit $ Credit $
A. Raw and in process inventory                   2,78,800
              account payable(6800×41)            2,78,800
B. Raw and in process inventory                   1,23,840
             conversion cost(6450×19.20)            1,23,840
C. Finished goods inventory(6320×(19.20+41))                   3,80,464
            raw and in process inventory            3,80,464
D. Sale account receivable                 11,76,960
            sales(6130×192)          11,76,960
D. Cost cost of goods sold(6130×(19.20+41))                   3,69,026
            finished goods inventory            3,69,026
4)
Ending balance of Raw and in process inventory:
$278,800+123,840 - $380,464 = $22,176
Finished goods inventory = Cost of goods manufactured - Cost of goods sold
= $380,464 - $3,69,026
   = $11,438